Shift Summary, Day shift 8/15/99

Yee Bob Hsiung (hsiung@fnal.gov)
Sun, 15 Aug 1999 15:34:14 -0500 (CDT)

Shift Summary for day shift, 15/Aug/99

On Shift: Sami Amasha, BobH

Hours of beam available at all intensities: 8.5h
Hours of beam available at nominal intensity: 8.25h
Hours of E832 data (special runs not included): 8h
Hours of special runs (laser scans, muon runs...):0.0h

Accounting of lost time:

KFRENDS - 7 times
Tape changes and failure init new run (drt_fresh_start twice).

========================================================================

Runs ended this shift:

Run 14232 - 23.7M events - BEAM_832 (34K Vac00)
Run 14235 - in progress.

Tapes brought to FCC this shift:
(including tapes for online split)

KQB093 QKK264-266 QKS264-266 QKC264-266
========================================================================

Accesses during shift: None.

========================================================================

Overview of shift:

Fairly steady beam, though SEM goes up to ~7.2E12 for a few spills.
Clean up ktev3a/data/CsI after 91% full warning.
drt_fresh_start twice to get new run with new tapes going.
Onsplit tape mounts for QKNL21, QKCC12, QKN023.

========================================================================

Description of Shift:

run____spill______time_______comment

14232 0752 SDAQ button turned Red (swic db too big)
notified Ashkan to fix it.
14232 150 0840 ktevana error - kfrend.
0847 onsplit tape mount QKNL21.
0903 " " " QKCC12.
174 0913 ktevana error - kfrend (low livetime warning).
198 0943 " " "
1045 ktev3a disk 91% full warning,
remove CsI data files up to run 14154.
297 1158 ktevana error - kfrend.
1228 onsplit tape mount QKN023.
328 1240 trigger hung - kfrend.
14232 362 1324 tape full, run stopped,
tape change and init new run but ktev3.
had all cpu's button turned RED.
drt_fresh_start twice to get everything going.
14235 1-2 1400 kfrend twice at the beginning.
51 1507 DAQDEAD_SEC = 1.69 sec when we got 7.1E12.
55 1514 " 1.52 6.7E12.

========================================================================

Follow up Notes on Previous Shift

DAQDEAD_SEC out of range messages did occur twice in this shift
maybe due to instantaneous rate at high intensity ~7E12.

========================================================================

Items to Follow up on Next Shift

No muon run and no magnet polarity changes till Monday Day shift.